OpenStudy (anonymous):

Derivative of Y=(x-1)/x ?

OpenStudy (anonymous):

(x-1)/x=1-1/x derivative=0-(-1/x^2)=1/x^2

OpenStudy (anonymous):

thanks! *bows*